OpenStudy (anonymous):

Concerning translation matrices..... How do you find a translation matrix to begin with??

OpenStudy (anonymous):

Let's make sure we are talking about the same thing. Like, say, translate a line right and down, that sort of thing?

OpenStudy (anonymous):

Yeah. The picture I'm supposed to be doing it for is this

OpenStudy (anonymous):

OK, translate the triangle somewhere?

OpenStudy (anonymous):

Pretty much

OpenStudy (anonymous):

Say you have a point (x,y), to translate it with t(a,b) you just put (x+a,y+b) So point A (-1-2), say you want to translate t(2,2) (2 up and 2 right) is (-1 +2 ,- 2 +2) You can write it in matrix notation if you want to

OpenStudy (anonymous):

there is more here http://en.wikipedia.org/wiki/Translation_matrix

OpenStudy (anonymous):

Well, it has to look like this;

OpenStudy (turingtest):

I don't see a B' or C' on your picture, so that makes no sense...

OpenStudy (turingtest):

are we to assume they all change in the same way A translates to A' ??

OpenStudy (anonymous):

It's just overkill, every point is being specified even with repetition.....

OpenStudy (anonymous):

It's just t(a,b) written 3 times....

OpenStudy (turingtest):

I don't see your meaning @estudier i see A, A', B, C, D' and E'

OpenStudy (anonymous):

It's a straight linear transformation t(a,b)

OpenStudy (turingtest):

left 3 and up 2 you mean, right?

OpenStudy (anonymous):

He just wants to write the transformation in matrices (that's what they asked him to do)

OpenStudy (turingtest):

x-3, y+2

OpenStudy (anonymous):

Yes, we've been there, done that....

OpenStudy (turingtest):

I see that

OpenStudy (anonymous):

We don't know what the actual transform is ....

OpenStudy (anonymous):

Can I go now?

OpenStudy (turingtest):

well that's what I'm saying, I suppose we must assume that for every point so what is the transform\[A\binom xy=\binom{x-3}{x+2}\]is the question, right?

OpenStudy (turingtest):

no @estudier you must stay or I will ban you forever :P nah, what's keeping you here?, I'll try it out

OpenStudy (anonymous):

y+2

OpenStudy (anonymous):

My computer's about to die and we lost power from this storm so I'm going to have to get back you in a little while

OpenStudy (turingtest):

my bad...

OpenStudy (turingtest):

then to be continued...

OpenStudy (anonymous):

Yes, to be continued

OpenStudy (anonymous):

Your teacher has made this unnecessarily difficult.....

OpenStudy (turingtest):

I agree

OpenStudy (anonymous):

I don't have a teacher. That's the problem

OpenStudy (anonymous):

Ah!

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

I'll be back on later guys

OpenStudy (anonymous):

Ok, it is actually very easy once you see it,......

OpenStudy (anonymous):

I'm back!

OpenStudy (anonymous):

I guess what you need is represent translation by multiplication? Tell me if I am right.

OpenStudy (anonymous):

@cerabear

OpenStudy (anonymous):

I think so. I'm not entirely sure

OpenStudy (anonymous):

if that is so: imagin the point you trying to translate as a position vector p: 

OpenStudy (anonymous):

now, you whant to translte it by vector V(a,b) 

OpenStudy (anonymous):

then your matrix is: \[\left[\begin{matrix}1 & 0 & a\\0&1 &b\\ 0 & 0 & 1\end{matrix}\right]\left(\begin{matrix}x\\ y\\ 1\end{matrix}\right)\]

OpenStudy (anonymous):

Oh. Well it's all pretty straight forward now lol

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

just to make it clear: x and y that apear in the column matrix are the coordinates of mpoint P. So P(x,y)

OpenStudy (anonymous):

@estudier @TuringTest

OpenStudy (anonymous):

Okay, cool. Thanks!

OpenStudy (anonymous):

you wellcome